# Transition Path Sampling with Improved Off-Policy Training of Diffusion Path Samplers

### v4

@@ 4.4 Chignolin folding
Finally, we consider a challenging molecule, Chignolin, an artificial protein consisting of 10 amino acids (Honda et al., 2004), which folds into a $\beta$-hairpin structure by hydrogen bonds. We sample transition paths from the unfolded state (right white circle) to the folded state (left white circle) in Figure 6. We define the target meta-stable state =$\mathcal{B}=\{\bm{R}~{}|~{}\lVert\xi(\bm{R})-\xi(\bm{R}_{\mathcal{B}})\rVert<0.75\}$ where $\xi$ consists of the top two components of time-lagged independent component analysis (TICA; Pérez-Hernández et al., 2013). We further describe TICA in Section B.4.
[…]
@@ 4.5 Ablation study
Effectiveness of algorithmic components. First, we conduct ablation studies to verify the effectiveness of the five proposed components: log-variance loss, learnable control variate, replay buffer, simulated annealing, and maximum over RBF values for various path lengths. To be specific, we (1) replace our loss with the KL divergence, (2) replace the learnable control variate with the local control variate used in Nüsken & Richter (2021), (3) remove the replay buffer and use data only from the current policy, (4) use only one temperature $\lambda$, and (5) remove maximum operation over RBF values using only the final state.
[…]
As seen in Figure 9, all the proposed components improve performance. Our loss is smaller than the KL divergence by more than two orders of magnitude and significantly improves performance. Learning the control variate slightly improves performance, showing that utilizing data from previous policies is effective. The replay buffer significantly improves training efficiency, and shows that the large performance gap between our loss and KL divergence comes from the replay buffer. Simulated annealing for biased MD simulation is critical to finding transition paths. RMSD does not decrease without simulated annealing while loss decreases significantly. For the relaxed indicator function, maximum operation accelerates convergence and improves performance with frequent training signals from the subtrajectories.
@@
(b) Loss and RMSD curve over rollouts.
Figure 9: Ablation studies on the components of TPS-DPS in the Alanine Dipeptide. (9(a)) Benchmark scores on Alanine Dipeptide. All metrics are averaged over 64 paths, and we highlight the best results in bold. ETS is computed for paths that hit the target meta-stable state. (9(b)) Loss and RMSD curves averaged over 8 different seeds.
Comparison to steered MD with varying force constant $k$. The performance of SMD differs by the force constant $k$. As seen in Figure 8, one can see the trade-off between THP and ETS for various $k$ values. TPS-DPS outperforms SMD in THP and ETS without showing a trade-off between them.
@@ 5 Conclusion
In this work, we introduced a novel CV-free diffusion path sampler, called TPS-DPS, to amortize transition path sampling, using log-variance divergence with the learnable control variate, and the replay buffer with simulated annealing for off-policy training. We also propose new $SE(3)$ equivariant scale-based parameterization for bias force and relaxed indicator function for frequent training signals. Evaluations on synthetic double-well and real-molecule systems such as Alanine Dipeptide, Polyproline Helix, and Chignolin demonstrate superior accuracy and diversity of our model compared to both classical and ML approaches.
Limitation.
While our algorithm successfully samples transition paths for small proteins, we did not evaluate our algorithm for even larger proteins due to the lack of computational budget. In addition, our work does not generalize across unseen pairs of meta-stable states or different molecular systems. These points to an interesting venue for future research, which would be more appealing for practical applications in drug discovery or material design.
@@ A.1 Log variance formulation
$\frac{\mathrm{d}\mathbb{P}_{\bm{v}_{\theta}}}{\mathrm{d}\mathbb{P}_{0}}(\bm{X})=\exp\left(\int^{T}_{0}(\bm{v}_{\theta}^{T}\Sigma^{-1})(\bm{X}_{t})\cdot\mathrm{d}\bm{X}_{t}-\int^{T}_{0}(\Sigma^{-1}\bm{u}\cdot\bm{v}_{\theta})(\bm{X}_{t})\mathrm{d}t-\frac{1}{2}\int^{T}_{0}\lVert\bm{v}_{\theta}(\bm{X}_{t})\rVert^{2}\mathrm{d}t\right)$ (14)
@@ A.2 Connection to existing loss functions on discrete-time domain
A.3 Training algorithm
Algorithm 1 Training
2:  for $i=1,\ldots,I$ do
5:     for $j=1,\ldots,J$ do
6:        Sample $K$ data $\{\bm{x}_{0:L}^{(k)}\}^{K}_{k=1}$ from $\hat{\mathcal{D}}$.
7:        Update $\theta$ and $w$ with the gradient of $\frac{1}{K}\sum_{k=1}^{K}\left(\log\frac{p_{0}(\bm{x}^{(k)}_{0:L})1_{\mathcal{B}}(\bm{x}^{(k)}_{0:L})}{p_{\bm{v}_{\theta}}(\bm{x}^{(k)}_{0:L})}-w\right)^{2}$
8:     end for
9:  end for
@@ B.1 OpenMM configurations
For real-world molecules, Alanine Dipeptide, Polyproline Helix, and Chingolin, we use the VVVR integrator (Sivak et al., 2014) with the step size $\Delta t=1~{}\mathrm{fs}$ and the friction term $\gamma=1~{}\mathrm{ps}^{-1}$. In the TPS-DPS training algorithm, we start simulations at a temperature $\lambda_{\text{start}}=600$K, and end at a temperature $\lambda_{\text{end}}=300$K. Other OpenMM configurations are shown in Table 2.
Table 2: OpenMM configurations for three real-world molecular systems.
Molecule Simulation time ($T$) Force field Solvent
Alanine Dipeptide $1000$ amber99sbildn (Lindorff-Larsen et al., 2010) vaccum
Polyproline Helix $10000$ amber/protein.ff14SBonlysc (Maier et al., 2015) implicit/gbn2
Chignolin $5000$ amber/protein.ff14SBonlysc implicit/gbn2
@@ B.2 Model configurations
We use a 3-layer MLP for the double-well system, and a 6-layer MLP for real-world molecules with ReLU activation functions for neural bias force, potential, and scale. To constrain the output of the neural bias scale parameterization to a positive value, we apply Softplus (Zheng et al., 2015) to the MLP output. As an input to the neural network, we concatenate the current position $(\bm{R}_{t})_{i}$ of the $i$-th atom with its distance to the target position $d_{i}=\lVert(\tilde{\bm{R}}_{\mathcal{B}})_{i}-(\bm{R}_{t})_{i}\rVert_{2}$. For real-world molecules, we apply the Kabsch algorithm (Kabsch, 1976) for heavy atoms to align $\bm{R}_{\mathcal{B}}$ with $\bm{R}_{t}$. We update the parameters of the neural network with a learning rate of $0.0001$, while the scalar parameter $w$ is updated with a learning rate of $0.001$. We clip the gradient norm with 1 to prevent loss from exploding. we train $J=1000$ times per rollout. We report other model configurations in Table 3. For PIPS, we use the model configurations reported by Holdijk et al. (2024). For CVs of SMD, we use backbone dihedral angles $(\phi,\psi)$ for Alanine Dipeptide and RMSD for Polyproline Helix and Chignolin.
@@ B.3 Evaluation metrics
Root mean square distance (RMSD).
We use the Kabsch algorithm (Kabsch, 1976) for heavy atoms to align the final position with the target position $\bm{R}_{\mathcal{B}}$, using the optimal (proper) rotation and translation to superimpose two heavy atom positions. We calculate RMSD between heavy atoms of the final position and the target position $\bm{R}_{\mathcal{B}}$.
Target hit percentage (THP).
THP measures the success rate of paths arriving at the target meta-stable state $\mathcal{B}$ in a binary manner. Formally, given the final positions $\{R^{(i)}\}^{M}_{i=1}$ of $M$ paths, THP is defined as follows:
@@
Energy of transition state (ETS).
ETS measures the ability of the method to find probable transition states when crossing the energy barrier. ETS refers to the maximum potential energy among states in a transition path. Formally, given a transition path $\bm{x}_{0:L}$ of length $L$ that reaches the target meta-stable state i.e., $\bm{R}_{L}\in\mathcal{B}$, ETS is defined as follows:
@@ B.4 System details
Double-well potential and dynamics.
Double-well system follows the overdamped Langevin dynamics defined as follows:
@@
Handedness.
For four points $A,B,C,D$, we define a handedness (Moradi et al., 2009) as follows:
$H_{ABCD}=\frac{\overrightarrow{EF}\cdot(\overrightarrow{CD}\times\overrightarrow{AB})}{|\overrightarrow{EF}|\cdot|\overrightarrow{CD}|\cdot|\overrightarrow{AB}|}.$ (37)
Here, $E$ and $F$ are the midpoints of the vectors $\overrightarrow{AB}$ and $\overrightarrow{CD}$, respectively. With the backbone atoms $X_{1},X_{2},...,X_{N}$ of a Polyproline Helix, we can define the handedness for $N$ atoms as $H=\sum_{i=1}^{N-3}H_{X_{i}X_{i+1}X_{i+2}X_{i+3}}$. In our experiments, we take the alpha carbon $X_{1}$, the carbonyl carbon $X_{2}$ of the first proline residue, the alpha carbon $X_{3}$ of the second residue and alpha carbon $X_{4}$ of third residues to compute the handedness.
Time-lagged independent components (TICA).
@@
Reproducibility.
We describe experiment details in Appendix B, including detailed simulation configuration and hyper-parameters. In the anonymous link, we provide the code for TPS-DPS.

### v5

@@ 4.3 Fast-folding Proteins
Finally, for challenging molecules, we consider four fast-folding proteins (Lindorff-Larsen et al., 2011): Chignolin, Trp-cage, BBA, and BBL with 10, 20, 28, and 47 amino acids, respectively. Proteins fold into stable structures by forming networks of hydrogen bonds. We aim to sample protein folding processes as seen in Figure 7. We define the target meta-stable state as =$\mathcal{B}=\{\bm{R}~{}|~{}\lVert\xi(\bm{R})-\xi(\bm{R}_{\mathcal{B}})\rVert<0.75\}$ where $\xi$ consists of the top two components of time-lagged independent component analysis (TICA; Pérez-Hernández et al., 2013). We further describe TICA in Section B.4.
As shown in Table 2 and Figure 7, only TPS-DPS (S) successfully samples transition paths that pass probable transition states while TPS-DPS (F) and TPS-DPS (P) fail to hit the target meta-stable state due to the lack of meaningful training signal. While SMD hits the target meta-stable, its transition paths pass less probable transition states. UMD and PIPS fail to hit the target meta-stable state. In Figure 5, we validate the sampled paths using the potential energy and donor-acceptor distance of the two key hydrogen bonds in Chignolin folding, ASP3OD-THR6OG and ASP3N-THR8O used in (Yang et al., 2024). The sampled path of reduces the donor-acceptor distance below the threshold $3.5$Å. For 3D videos of transition paths of fast-folding proteins, we refer to project page.
@@ 4.4 Ablation study
We conduct ablation studies to verify the effectiveness of the six proposed components: log-variance loss, learnable control variate, replay buffer, simulated annealing, maximum over RBF values for various path lengths, and $SE(3)$ equivariance. To be specific, we (1) replace the log-variance divergence with the KL divergence, (2) replace the learnable control variate $w\in\mathbb{R}$ with the local control variate which is Monte-Carlo estimator used in Nüsken & Richter (2021), (3) remove the replay buffer and rely solely on data from the current policy, (4) use only one temperature $\lambda=300$K in sampling, (5) remove maximum operation over RBF kernel values using only the final state, (6) skip the alignment of neural network input states with the target position by the Kabsch algorithm.
As seen in Figure 6, all the proposed components improve performance. Our loss is smaller than the KL divergence by more than two orders of magnitude and significantly improves performance. Learning the control variate slightly improves performance, showing that utilizing data from previous policies is effective. The replay buffer significantly improves training efficiency, and shows that the large performance gap between our loss and KL divergence comes from the replay buffer. Simulated annealing for biased MD simulation is critical to finding transition paths. RMSD does not decrease without simulated annealing while loss decreases significantly. For the relaxed indicator function, maximum operation accelerates convergence and improves performance with frequent training signals from the subtrajectories. Leveraging the symmetry of the bias force with the Kabsch algorithm improves performance. We further compare with reverse KL divergence in Appendix E.
@@
(b) Loss and RMSD curves over rollouts.
Figure 6: Ablation studies on each proposed component of TPS-DPS (F) in Alanine Dipeptide. (6(a)) Benchmark scores on Alanine Dipeptide. (6(b)) Loss and RMSD curves averaged over 8 seeds.
@@ 5 Conclusion
In this work, we introduced a novel CV-free diffusion path sampler, called TPS-DPS, to amortize the cost of sampling transition paths. We propose the log-variance divergence with the learnable control variate and off-policy training with the replay buffer and simulated annealing. We propose a new scale-based equivariant parameterization of bias force and relaxed indicator function for reaching target meta-stable states and frequent training signals, particularly in large molecules. Evaluations on double-well, Alanine Dipeptide, and four fast-folding proteins demonstrate that TPS-DPS is superior in accuracy and diversity compared to non-ML and ML approaches.
Limitation. While our experiments show promise, they are limited to small fast-folding proteins (up to 50 amino acids). The applicability of our method to real-world proteins with more than 500 amino acids remains unexplored. Additionally, integrating various libraries, such as PLUMED (plu, 2019) and DMFF (Wang et al., 2023), for MD with neural network bias force, has not yet been investigated.
Our method does not generalize across unseen pairs of meta-stable states or different molecular systems. These points to an interesting venue for future research, which would be more appealing for practical applications in drug discovery or material design.
@@ A.1 Log variance formulation
$\frac{\mathrm{d}\mathbb{P}_{\bm{v}_{\theta}}}{\mathrm{d}\mathbb{P}_{0}}(\bm{X})=\exp\left(\int^{T}_{0}(\bm{v}_{\theta}^{T}\Sigma^{-1})(\bm{X}_{t})\cdot\mathrm{d}\bm{X}_{t}-\int^{T}_{0}(\Sigma^{-1}\bm{u}\cdot\bm{v}_{\theta})(\bm{X}_{t})\mathrm{d}t-\frac{1}{2}\int^{T}_{0}\lVert\bm{v}_{\theta}(\bm{X}_{t})\rVert^{2}\mathrm{d}t\right).$ (14)
@@ A.2 Connection to existing loss functions on discrete-time domain
A.3 Proof of scale-based parameterization
In this section, we prove that our scale-based parameterization of bias force strictly decreases the distance to the (aligned) target position for small step sizes, improving the ability to find informative paths in large molecules.
Proposition 1.
Consider the molecular state $\bm{R}_{t}$ at the $t$-th time step and the next state $\bm{R}_{t+\Delta t}^{\prime}=\bm{R}_{t}+\bm{b}(\bm{X}_{t})\Delta t/\bm{m}$ updated by step size $\Delta t$ and the bias force $\bm{b}(\bm{X}_{t})=\text{diag}(\bm{s}_{\theta}(\rho_{t}^{-1}\cdot\bm{X}_{t}))(\rho_{t}\cdot{\bm{R}}_{\mathcal{B}}-\bm{R}_{t})$. Then there always exists a small enough $\Delta t$ that strictly decreases the distance towards the target state $\bm{R}_{\mathcal{B}}$:
$\lVert\rho^{\prime}_{t+\Delta t}\cdot\bm{R}_{\mathcal{B}}-\bm{R}^{\prime}_{t+\Delta t}\rVert<\lVert\rho_{t}\cdot\bm{R}_{\mathcal{B}}-\bm{R}_{t}\rVert,$ (33)
where $\rho^{\prime}_{t+\Delta t}=\operatorname{argmin}_{\rho\in SE(3)}\lVert\rho\cdot\bm{R}_{\mathcal{B}}-\bm{R}^{\prime}_{t+\Delta t}\rVert$ and we assume that there does not exist a rotation that exactly aligns the current molecular state to the target state, i.e., $\lVert\rho_{t}\cdot\bm{R}_{\mathcal{B}}-\bm{R}_{t}\rVert>0$.
Proof.
The proof consists of two steps. We first show the (strictly) positive correlation between the bias force and the direction from the $t$-th state $\bm{R}_{t}$ to the target state $\bm{R}_{\mathcal{B}}$. Next, we show that the positive correlation gaurantees a strict decrease in distance between the states, i.e., $\lVert\rho_{t}\cdot\bm{R}_{\mathcal{B}}-\bm{R}_{t}\rVert$, given that the distance was not already zero.
Step 1: First, we show that the bias force (divided by atom-wise masses) is positively correlated with the direction to the target position, i.e., $(\bm{b}(\bm{X}_{t})/\bm{m})^{\top}(\rho_{t}\cdot\bm{R}_{\mathcal{B}}-\bm{R}_{t})>0.$ This follows from:
$(\bm{b}(\bm{X}_{t})/\bm{m})^{\top}(\rho_{t}\cdot\bm{R}_{\mathcal{B}}-\bm{R}_{t})$ $=(\rho_{t}\cdot\bm{R}_{\mathcal{B}}-\bm{R}_{t})^{T}\frac{\text{diag}(\bm{s}_{\theta}(\rho_{t}^{-1}\cdot\bm{X}_{t}))}{\bm{m}}(\rho_{t}\cdot\bm{R}_{\mathcal{B}}-\bm{R}_{t})$ (34)
$=\sum_{i=1}^{3N}\left(\frac{\bm{s}_{i}}{\bm{m}_{i}}\right)\left(\rho_{t}\cdot\bm{R}_{\mathcal{B}}-\bm{R}_{t}\right)^{2}_{i}>0,$ (35)
where $\bm{s}_{i}>0$ is the $i$-th element of $\bm{s}_{\theta}(\rho_{t}^{-1}\cdot\bm{X}_{t})$ and $\left(\rho_{t}\cdot\bm{R}_{\mathcal{B}}-\bm{R}_{t}\right)_{i}$ is the $i$-th element of the direction to the target position.
Step 2: Next, we show that the positive correlation ensures distance reduction for a small enough step size. Consider the squared distance between the target position $\rho_{t}\cdot\bm{R}_{\mathcal{B}}$ and updated position $\bm{R}^{\prime}$ by bias force
$\lVert\rho_{t}\cdot\bm{R}_{\mathcal{B}}-\bm{R}^{\prime}_{t+\Delta t}\rVert^{2}$
$=\lVert\rho_{t}\cdot\bm{R}_{\mathcal{B}}-(\bm{R}_{t}+\bm{b}(\bm{X}_{t})\Delta t/\bm{m})\rVert^{2}$ (36)
$=\lVert(\rho_{t}\cdot\bm{R}_{\mathcal{B}}-\bm{R}_{t})-\bm{b}(\bm{X}_{t})\Delta t/\bm{m}\rVert^{2}$ (37)
$=\lVert\rho_{t}\cdot\bm{R}_{\mathcal{B}}-\bm{R}_{t}\rVert^{2}-2\Delta t(\bm{b}(\bm{X}_{t})/\bm{m})^{\top}(\rho_{t}\cdot\bm{R}_{\mathcal{B}}-\bm{R}_{t})+(\Delta t)^{2}\lVert\bm{b}(\bm{X}_{t})/\bm{m}\rVert^{2}.$ (38)
Due to step 1, i.e., $(\bm{b}(\bm{X}_{t})/\bm{m})^{\top}(\rho_{t}\cdot\bm{R}_{\mathcal{B}}-\bm{R}_{t})>0$, there exists a step size $\Delta t$ satisfying:
$0<\Delta t<\frac{2(\bm{b}(\bm{X}_{t})/\bm{m})^{\top}(\rho_{t}\cdot\bm{R}_{\mathcal{B}}-\bm{R}_{t})}{\lVert\bm{b}(\bm{X}_{t})/\bm{m}\rVert^{2}}.$ (39)
With this choice of $\Delta t$, multiplying $\Delta t\lVert\bm{b}(\bm{X}_{t})/\bm{m}\rVert^{2}$ leads to the following inequaliity:
$(\Delta t)^{2}\lVert\bm{b}(\bm{X}_{t})/\bm{m}\rVert^{2}<2\Delta t(\bm{b}(\bm{X}_{t})/\bm{m})^{\top}(\rho_{t}\cdot\bm{R}_{\mathcal{B}}-\bm{R}_{t}).$ (40)
By subtracting the right-hand side from both sides and adding $\lVert\rho_{t}\cdot\bm{R}_{\mathcal{B}}-\bm{R}_{t}\rVert^{2}$ to both sides, we have the following inequality:
$\lVert\rho_{t}\cdot\bm{R}_{\mathcal{B}}-\bm{R}^{\prime}_{t+\Delta t}\rVert^{2}<\lVert\rho_{t}\cdot\bm{R}_{\mathcal{B}}-\bm{R}_{t}\rVert^{2}.$ (41)
Taking the square root of both sides, we have the following inequality:
$\lVert\rho^{\prime}_{t+\Delta t}\cdot\bm{R}_{\mathcal{B}}-\bm{R}^{\prime}_{t+\Delta t}\rVert\leq\lVert\rho_{t}\cdot\bm{R}_{\mathcal{B}}-\bm{R}^{\prime}_{t+\Delta t}\rVert<\lVert\rho_{t}\cdot\bm{R}_{\mathcal{B}}-\bm{R}_{t}\rVert,$ (42)
where the first inequality follows from the definition of $\rho^{\prime}_{t+\Delta t}=\operatorname{argmin}_{\rho\in SE(3)}\lVert\rho\cdot\bm{R}_{\mathcal{B}}-\bm{R}^{\prime}_{t+\Delta t}\rVert$. This completes the proof. ∎
@@ B.1 OpenMM configurations
For real-world molecules, we use the VVVR integrator (Sivak et al., 2014) with the step size $\Delta t=1~{}\mathrm{fs}$ and the friction term $\gamma=1~{}\mathrm{ps}^{-1}$. In the TPS-DPS training algorithm, we simulate MD with $T=1$ps for Alanine Dipeptide and $T=5$ps for the fast-folding proteins. We start simulations at a temperature $\lambda_{\text{start}}=600$K, and end at a temperature $\lambda_{\text{end}}=300$K for Alanine Dipeptide and Chignolin and $\lambda_{\text{end}}=400$K for the others. We use the amber99sbildn (Lindorff-Larsen et al., 2010) force field for Alanine Dipeptide in vacuum and the ff14SBonlysc (Maier et al., 2015) force field for the fast-folding proteins with the gbn2 implicit solvation model (Nguyen et al., 2013).
@@ B.2 Model configurations
We use a 3-layer MLP for the double-well system, and a 6-layer MLP for real-world molecules with ReLU activation functions for neural bias force, potential, and scale. To constrain the output of the neural bias scale parameterization to a positive value, we apply Softplus to the MLP output. As an input to the neural network, we concatenate the current position $(\bm{R}_{t})_{i}$ of the $i$-th atom with its distance to the target position $d_{i}=\lVert(\tilde{\bm{R}}_{\mathcal{B}})_{i}-(\bm{R}_{t})_{i}\rVert_{2}$. For real-world molecules, we apply the Kabsch algorithm (Kabsch, 1976) for heavy atoms to align $\bm{R}_{\mathcal{B}}$ with $\bm{R}_{t}$. We update the parameters of the neural network with a learning rate of $0.0001$, while the scalar parameter $w$ is updated with a learning rate of $0.001$. We clip the gradient norm with 1 to prevent loss from exploding. we train $J=1000$ times per rollout. We report other model configurations in Table 3. For PIPS, we use the model configurations reported by Holdijk et al. (2024). For CVs of SMD, we use backbone dihedral angles $(\phi,\psi)$ for Alanine Dipeptide and RMSD for fast-folding proteins.
@@ B.3 Evaluation metrics
Root mean square distance (RMSD). We use the Kabsch algorithm (Kabsch, 1976) for heavy atoms to align the final position with the target position $\bm{R}_{\mathcal{B}}$, using the optimal (proper) rotation and translation to superimpose two heavy atom positions. We calculate RMSD between heavy atoms of the final position and the target position $\bm{R}_{\mathcal{B}}$.
Target hit percentage (THP). THP measures the success rate of paths arriving at the target meta-stable state $\mathcal{B}$ in a binary manner. Formally, given the final positions $\{R^{(i)}\}^{M}_{i=1}$ of $M$ paths, THP is defined as follows:
@@
Energy of transition state (ETS). ETS measures the ability of the method to find probable transition states when crossing the energy barrier. ETS refers to the maximum potential energy among states in a transition path. Formally, given a transition path $\bm{x}_{0:L}$ of length $L$ that reaches the target meta-stable state i.e., $\bm{R}_{L}\in\mathcal{B}$, ETS is defined as follows:
@@ B.4 System details
Systhetic double-well system. Double-well system follows the overdamped Langevin dynamics defined as follows:
@@
Reproducibility. We describe experiment details in Appendix B, including detailed simulation configuration and hyper-parameters. In the anonymous link, we provide the code for TPS-DPS.
Appendix C Computational cost
In this section, we analyze the time complexity of TPS-DPS and provide the number of energy evaluations and runtime in training and inference time for real molecules.
The training and inference time complexity of TPS-DPS is $O(NMLJ)$ and $O(NML)$, respectively, where $N$ is the number of atoms, $M$ is the number of samples, $L$ is the number of MD steps, and $J$ is the number of rollouts. To be specific, training consists of biased MD simulations with $O(NML)$ time complexity. Given $M$, the total complexity of one biased MD step is $O(N)$.
To justify it, we note that the biased MD step consists of three stages: (1) calculating bias force, (2) calculating OpenMM force field, and (3) integrating the biased MD. Given the number of layers, and hidden units, MLP for bias force requires $O(N)$ and the Kabsch algorithm for equivariance requires $O(N)$. Calculating force with cut-off and integrating MD with VVVR integrator also requires $O(N)$.
To measure computational cost, we consider the number of energy evaluations and runtime per rollout in training and inference time. As shown in Table 4, the inference cost of TPS-DPS is proportional to UMD and SMD which have time complexity $O(NML)$. TPS-DPS requires less energy evaluations than PIPS in training since TPS-DPS finds transition paths faster than PIPS by utilizing the replay buffer and simulated annealing.
Table 4: Cost comparison. EET and EEI refer to the number of energy evaluations by the OpenMM library in training and inference, respectively. RT and RI denote runtime (second) per rollout in training and inference on a single RTX A5000 GPU. MD simulations are conducted with $T=1$ps for Alanine Dipeptide, $T=5$ps for other systems, and $\Delta t=1$fs. The number of samples for each training rollout and inference is 16 and 64, respectively. Other configurations of TPS-DPS are provided in Section B.2 and those of PIPS can be found in (Holdijk et al., 2024)
Molecule Method EET ($\downarrow$) EEI ($\downarrow$) RT ($\downarrow$) RI ($\downarrow$)
Alanine Dipeptide UMD - 64K - 29.49
SMD - 64K - 47.45
PIPS (F) 240M 64K 44.22 71.05
PIPS (P) 240M 64K 50.54 75.67
TPS-DPS (F, Ours) 16M 64K 24.93 70.50
TPS-DPS (P, Ours) 16M 64K 27.25 78.83
TPS-DPS (S, Ours) 16M 64K 25.11 73.04
Chignolin UMD - 320K - 224.23
SMD - 320K - 380.37
PIPS (F) 40M 320K 277.29 565.58
PIPS (P) 40M 320K 317.08 622.87
TPS-DPS (F, Ours) 8M 320K 209.29 562.90
TPS-DPS (P, Ours) 8M 320K 224.36 623.63
TPS-DPS (S, Ours) 8M 320K 215.18 581.26
Trp-Cage UMD - 320K - 258.29
SMD - 320K - 323.52
PIPS (F) 40M 320K 360.55 652.59
PIPS (P) 40M 320K 417.93 718.05
TPS-DPS (F, Ours) 8M 320K 289.10 655.22
TPS-DPS (P, Ours) 8M 320K 301.76 699.44
TPS-DPS (S, Ours) 8M 320K 293.51 673.00
BBA UMD - 320K - 395.12
SMD - 320K - 551.04
PIPS (F) 40M 320K 506.57 1080.62
PIPS (P) 40M 320K 574.07 1117.74
TPS-DPS (F, Ours) 8M 320K 422.23 1042.81
TPS-DPS (P, Ours) 8M 320K 430.24 1091.97
TPS-DPS (S, Ours) 8M 320K 426.48 1068.68
BBL UMD - 320K - 673.55
SMD - 320K - 853.77
PIPS (F) 40M 320K 622.22 1558.53
PIPS (P) 40M 320K 660.02 1612.78
TPS-DPS (F, Ours) 8M 320K 560.95 1520.05
TPS-DPS (P, Ours) 8M 320K 572.77 1607.62
TPS-DPS (S, Ours) 8M 320K 563.45 1553.89
Appendix D Visualization of sampled paths of fast folding proteins
(a) UMD
(b) SMD (20k)
(c) PIPS (P)
(d) TPS-DPS (F)
(e) TPS-DPS (P)
(f) TPS-DPS (S)
Figure 7: Visualization of sampled paths of four fast-folding proteins on free energy landscapes for top two TICA components. We aim to sample folding processes for the four fast folding proteins: Chignolin, Trpcage, BBA, and BBL (from top to bottom rows).
Appendix E Comparision with reverse KL divergence
Table 5: Benchmark scores of reverse KL divergence and TPS-DPS on Alanine Dipeptide system. Metrics are averaged over 64 paths. TPS-DPS consistently outperforms reverse KL divergence on all metrics regardless of predicting bias force or potential.
Method RMSD ($\downarrow$) THP ($\uparrow$) ETS ($\downarrow$)
Å % $\mathrm{kJmol}^{-1}$
Reverse KL (F) 0.43 $\pm$ 0.34 53.12 27.88 $\pm$ 14.38
Reverse KL (P) 0.58 $\pm$ 0.34 48.43 21.61 $\pm$ 11.76
TPS-DPS (F, Ours) 0.16 $\pm$ 0.06 92.19 19.82 $\pm$ 15.88
TPS-DPS (P, Ours) 0.16 $\pm$ 0.10 87.50 18.37 $\pm$ 10.86
(a) Reverse KL (F)
(b) Reverse KL (P)
(c) TPS-DPS (F)
(d) TPS-DPS (P)
Figure 8: Visualization of sampled paths of Alanine Dipeptide on the Ramachandran plot. The reverse KL divergence struggles to find diverse reaction channels suffering from mode collapse issues while the log-variance divergence of our method can capture two reaction channels and reach the target states better.